OpenStudy (anonymous):

add. v/v^2-5v-24+6/v+3 the sum is...

OpenStudy (saifoo.khan):

please rewrite.

OpenStudy (anonymous):

^^

OpenStudy (anonymous):

vv^2-5v-24+6v+3

OpenStudy (anonymous):

that looks horrible

OpenStudy (saifoo.khan):

lol.

OpenStudy (anonymous):

its v over v^2-5v-24 plus 6 over v+3

OpenStudy (saifoo.khan):

\[\frac{v}{v^2-5v-24} + \frac{6}{v+3}\]

OpenStudy (anonymous):

yes!how did you do that

OpenStudy (saifoo.khan):

are u doing mathlab?

OpenStudy (anonymous):

ya.lol

OpenStudy (saifoo.khan):

lol, cool.

OpenStudy (anonymous):

ya, but i cant comprehend it

OpenStudy (saifoo.khan):

Yes, i can understand.. i know many people who can't understand algebra. :(

OpenStudy (saifoo.khan):

\[\frac{7 v-48}{v^2-5v-24}\]

OpenStudy (anonymous):

ya. it sucks. This is my last class and my instructor told me to quit 3 weeks ago

OpenStudy (saifoo.khan):

ohh.. which week are u in?

OpenStudy (anonymous):

5

OpenStudy (saifoo.khan):

wow, alot to go ahead. best of luck! :D

OpenStudy (anonymous):

thanks

OpenStudy (saifoo.khan):

np

OpenStudy (saifoo.khan):

i have actually done these questions many times.. so i remember them.. lol.

OpenStudy (anonymous):

oh awesome!

OpenStudy (saifoo.khan):

hehe. yea/